What is the approximate ratio of lactic acid molecules to lactate ions in a urine sample at pH 1.9? (lactic acid pK=3.9)

THE ANSWER IS 100:1 BUT I NEED A THOROUGH EXPLANATION I have no idea how to do this problem at all.
Answer #1

As we know that

pH = pKa + log {base / weak acid}

pH - pKa = log { base / weak acid }

1.9 - 3.9 = log { base / weak acid }

-2 = log { base / weak acid }

10^-2 = base / weak acid

1/100 = base / weak acid

So the ratio is such as

Weak acid / base = lactic acid / lactate ion = 100/1

So ratio is 100:1